what is the center of the circle defined y the equation x^2+2x+y^2=15
Answer by lwsshak3(11628)

what is the center of the circle defined y the equation x^2+2x+y^2=15
***
Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=coordinates of center, r=radius
..
x^2+2x+y^2=15
complete the square:
(x^2+2x+1+y^2=15+1
(x+1)^2+y^2=16
center:(-1,0)